OpenStudy (anonymous):

Copy the graph of y=f(x) and draw the graphs of y=-f(x) and y=f(-x) on the same axes. HELP!

OpenStudy (anonymous):



OpenStudy (anonymous):

that is the graph of y=f(x) and I don't know how to draw the graphs of y=-f(x) and y=(-x)..

OpenStudy (apoorvk):

y=f(-x) is graph of the function about negative half of the x-axis. y=-f(x) is the flipped out graph of the whole function about about the whole x-axis (basically mirror image)

OpenStudy (apoorvk):

try and post and let us check it.

OpenStudy (anonymous):

 like that for y=(-x)?